# SurgiTrack: Fine-Grained Multi-Class Multi-Tool Tracking in Surgical Videos

### v1

@@ 6.9 Tracking Evaluation under Surgical Visual Challenges
Conversely, in crowded scenes, the model faces challenges as numerous tools closely packed together hinder detection accuracy, potentially affecting accurate tracking. Disentangling cluttered tools proves difficult, potentially affecting accurate identity matching (-4.1% AssA, -4.4% AssRe). Smoke from coagulation instruments decreases the tracking score due to impaired visibility, resulting in significant missed detections (-9.9% DetA, -11.1% DetRe) and diminishing association accuracy.
@@
Limited camera coverage poses a major challenge, causing numerous missed detections, although localizing focused tools remains precise. Constrained trocar view hampers the model’s ability to track and manage tool identities, especially complicating the interpretation of tool direction, as a link to the operating surgeon’s hand. Specular light reflection emerges as the most impactful visual challenge, leading to a notable decrease in all tracking aspects, with a huge loss of 13.5% points in HOTA. The excessive brightness from specular light obscures visual details, leading to a 16.8% and 9.9% reduction in detection and association accuracies respectively.
@@ Acknowledgements
This work was supported by French state funds managed within the Plan Investissements d’Avenir by the ANR under references: National AI Chair AI4ORSafety [ANR-20-CHIA-0029-01], IHU Strasbourg [ANR-10-IAHU-02] and by BPI France [Project 5G-OR]. This work was granted access to the servers/HPC resources managed by CAMMA, IHU Strasbourg, Unistra Mesocentre, and GENCI-IDRIS [Grant 2021-AD011011638R3, 2021-AD011011638R4]. The authors acknowledge the suggestions of Rupak Bose on the visualization plots and thank Dr. Joël L. Lavanchy for some useful clinical discussions.

### v3

@@ 6.8 Tracking Evaluation under Surgical Visual Challenges
Conversely, in crowded scenes, the model faces challenges as numerous tools closely packed together hinder detection accuracy. Disentangling cluttered tools proves difficult, potentially affecting accurate identity matching (-4.1% AssA, -4.4% AssRe). Smoke from coagulation tools decreases the tracking score due to impaired visibility, resulting in significant missed detections (-9.9% DetA, -11.1% DetRe) and diminishing association accuracy.
@@
Limited camera coverage poses a major challenge, causing numerous missed detections, although localizing focused tools remains precise. Constrained trocar view hampers the model’s ability to track and manage tool identities, especially complicating the interpretation of tool direction, as a link to the operating surgeon’s hand. Specular light reflection emerges as the most impactful visual challenge, leading to a notable decrease in all tracking aspects, with a huge loss of 13.5% points in HOTA. The excessive brightness from specular light obscures visual details, leading to a 16.8% and 9.9% reduction in detection and association accuracies respectively. Though rare in this dataset, appearance changes in a single tool can also result from the use of electrocautery instruments.
@@
Beyond the failure modes due to the surgical visual challenges, we observe that the model experiences difficulties at the beginning of a video due to limited historical data for the HGBM to rely on, leading to more frequent identity fragmentations, however, this usually stabilizes in few seconds.
As a limitation, our approach would not detect a replacement of a tool by another tool of the same class (and same manufacturer) if operated via the same trocar port. This is because the proposed direction feature is designed to uniquely differentiate tools based on their originating directions - the trocar ports. Additional signals would be needed to tackle this situation. The use of direction features may face challenges in certain type of surgery such as endoscopic endonasal surgery where the nostrils are closely space, or where trocars are closely spaced with little or no directional difference.
@@ Acknowledgements
This work was supported by French state funds managed within the Plan Investissements d’Avenir by the ANR under references: National AI Chair AI4ORSafety [ANR-20-CHIA-0029-01], IHU Strasbourg [ANR-10-IAHU-02] and by BPI France [Project 5G-OR]. This work was granted access to the servers/HPC resources managed by CAMMA, IHU Strasbourg, Unistra Mesocentre, and GENCI-IDRIS [Grant 2021-AD011011638R3, 2021-AD011011638R4]. The authors acknowledge the suggestions of Engr. Rupak Bose on the visualization plots and thank Dr. Joël L. Lavanchy for some useful clinical discussions.